yesno:

hehe.... i need a lot of help.. Golf balls come in a variety of packages. A particular brand of golf ball comes in a package of 3 golf balls and a package of 12 golf balls. The package with 3 golf balls costs $3.36, and the package with 12 golf balls costs $6.90. Which is the better deal? What is the unit price of the better deal? a. 12-ball pack at $1.12 per golf ball b. 3-ball pack at $0.58 per golf ball c. 3-ball pack at $0.28 per golf ball d. 12-ball pack at $0.58 per golf ball

jimthompson5910:

unit price = (total cost)/(number of items) Let's say you pay a total of $18 for 6 items, as an example The unit cost would be 18/6 = 3 dollars per item

Maddie2005:

idk about this now but maybe D

Maddie2005:

sorry if u get it wrong

yesno:

no its okay, thanks for trying to help :)

Maddie2005:

yw

yesno:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jimthompson5910 unit price = (total cost)/(number of items) Let's say you pay a total of $18 for 6 items, as an example The unit cost would be 18/6 = 3 dollars per item \(\color{#0cbb34}{\text{End of Quote}}\) i- i don't understand 😃👍🏽

jimthompson5910:

Let's say one orange is $2 If we buy 5 oranges, then 5*2 = 10 dollars is the total cost, agreed?

yesno:

agreed? I'm dumb bare w me.

jimthompson5910:

So what we do is just go in reverse of that We divide 10 over 5 10/5 = 2 If we spend $10 total and buy 5 oranges, then the unit cost is $2 In other words, the cost per orange is $2

yesno:

i still dont understand- um I'm very stupid.

Maddie2005:

A= 13.44 B= 1.74 C= 0.84 D=6.96

yesno:

I'm jus gon go w D. oh well if i get it wrong.

Maddie2005:

no i think it's c now

yesno:

okay then c-

Maddie2005:

yea go with C

yesno:

kk

CourtneyWatson:

The better deal is the 12 pack because the unit price is $0.37 which is lower than the unit price of the 18 pack that is $0.80.